# Generalized Uncertainty Principle theories and their classical interpretation

### v1

@@ 4.2 Kempf-Mangano-Mann case in Maggiore’s scheme
which is the correct expression of the KMM algebra at classical level and the same formula obtained in Section 3.3.2, with the identification previously mentioned, i.e. $S_{ij}=0$ and $c=1$.
5 Proof of the Theorem
In this Section, we are going to provide a proof for the Theorem 2.3. Let us start from the following Proposition (Prop.8.11 in [15]):
@@ 5 Proof of the Theorem
A symplectic form on $\mathcal{M}$ is equivalent to a nondegenerate Poisson structure.
@@
Consider the fundamental Poisson brackets expressed in local coordinates $x^{a}=(q_{1},\dots,q_{d},p_{1},\dots,p_{d})$ as in (2.1) , where $L_{ij}$ and $f$ are smooth functions of the coordinates, and $f$ is strictly positive on $\mathcal{M}$. If the Poisson brackets satisfy the Jacobi identities, then they define a nondegenerate Poisson structure. The non degeneracy can easily checked considering the bivector field in coordinate
@@
for which, $\det(\pi^{ab})=f^{2d}>0$. From this and Prop. 5.1, the following Lemma holds:
@@
Now, we want to prove that if the Poisson brackets satisfy the Jacobi identity in the classical interpretation of a GUP theory, then, the algebra of the commutators of the quantum theory also satisfy the Jacobi identity.
Since a rigorous and consistence quantization procedure does not exist for GUP theories, we just check that the quantum Jacobi identity are satisfied if we consider $f$ and $L$ as in Section 3.3 for two possible operator ordering.
@@ 6 Conclusions
Finally, we concluded the article with a demonstration that bridges the classical theory thus formulated and the quantum theory: a properly defined classical GUP theory can be correctly interpreted if and only if the corresponding quantum commutators satisfy the Jacobi identities. While transitioning from quantum to classical theory is straightforward concerning the Jacobi identities for brackets and commutators, the reverse is not obvious due to potential operator ordering ambiguities. If all classical relations ensuring a symplectic structure in GUP theory are assumed valid at the quantum level, this will ensure that the Jacobi identities for all commutators without ordering ambiguities are automatically satisfied. For the remaining commutators, specifically those involving only the generalized coordinates $q$s, verifying the Jacobi identity requires fixing an ordering. Although proving that any ordering satisfies the Jacobi identities is generally not possible, the two most natural orderings do, providing a guideline for constructing the quantum theory and hence concluding the demonstration.

### v2

@@ 4.2 Kempf-Mangano-Mann case in Maggiore’s scheme
which is the correct expression of the KMM algebra at the classical level and the same formula obtained in Section 3.3.2, with the identification previously mentioned, i.e. $S_{ij}=0$ and $c=1$.
5 Verifying the Conjecture
In this Section, we are going to provide some cases that satisfy the Conjecture 2.3. Let us start from the following Proposition (Prop.8.11 in [16]):
@@ 5 Verifying the Conjecture
A symplectic form on $\mathcal{M}$ is equivalent to a non-degenerate Poisson structure.
@@
Consider the fundamental Poisson brackets expressed in local coordinates $x^{a}=(q_{1},\dots,q_{d},p_{1},\dots,p_{d})$ as in (2.1) , where $L_{ij}$ and $f$ are smooth functions of the coordinates, and $f$ is strictly positive on $\mathcal{M}$. If these Poisson brackets satisfy the Jacobi identities, then they define a non-degenerate Poisson structure. The non-degeneracy can be easily checked considering the bivector field in coordinates
@@
for which $\det(\pi^{ab})=f^{2d}>0$. From this and Prop. 5.1, the following Lemma holds:
@@
Since a rigorous and consistent quantization procedure does not exist for GUP theories, we just check that the quantum Jacobi identity is satisfied if we consider $f$ and $L$ as in Section 3.3 for two possible ’natural’ operator ordering in the quantum theory, namely all the momenta to the left and all the momenta to the right.
@@ 6 Conclusions
An interesting follow-up to the analysis we carried out regarding the classical symplectic structure of these theories could be represented by the investigation of relativistic non-commutative models. In this respect, the well-known Snyder space-time [22] is a notable representative that will deserve examination within our framework.
Finally, we concluded the article by discussing a Conjecture in order to bridge the classical theory thus formulated and the quantum theory: a properly defined GUP theory can be correctly classically interpreted if and only if the corresponding quantum commutators satisfy the Jacobi identities. While transitioning from quantum to classical theory is straightforward concerning the Jacobi identities for brackets and commutators, the reverse is not obvious due to potential operator ordering ambiguities. If all classical relations ensuring a symplectic structure in GUP theory are assumed valid at the quantum level, this will ensure that the Jacobi identities for all commutators without ordering ambiguities are automatically satisfied. For the remaining commutators, specifically those involving only the generalized coordinates $q$s, verifying the Jacobi identity requires fixing an ordering. Although proving that any ordering satisfies the Jacobi identities is generally not possible, the two most natural orderings do, providing potentially a guideline for constructing the quantum theory.